In a 5.3 L of a 3.5 M solution of KCL there are _____ molds of KCL
1
Expert's answer
2018-06-06T07:31:58-0400
V=5.3 L
CM(KCl)=3.5 mol
.n (KCl) =?
.n=CM×V=3.5×5.3=18.55 mol